OpenStudy (anonymous):

Surface Area: 12 in 2 = Length 8 in 2 = Width 2 1/2 in 2 = Height

OpenStudy (anonymous):

@AmazingApples00 @bcimkenz @DiamondBear3299 @dinopyte @Here_to_Help15 @Michele_Laino @shifuyanli @sleepyjess @CausticSyndicalist

OpenStudy (anonymous):

@Michele_Laino can you help me?

OpenStudy (michele_laino):

If your solid shape is a parallelepiped, then please you have to perform this computation, in order to find the total area S: S=2*[(12*8)+(8*5/2)+(12*5/2)] since 2 1/2=2+1/2=5/2 and Length=12 inches, Width= 8 inches and height= 2+1/2 inches